Suppose 1.5 percent of the antennas on new Nokia cell phones are defective. For a random sample of 200 antennas, find the probability that:
1. None of the antennas is defective.
2. Three or more of the antennas are defective

Answer by stanbon(75887)

Suppose 1.5 percent of the antennas on new Nokia cell phones are defective.
For a random sample of 200 antennas, find the probability that:
---
Binomial Problem: n = 200 ; p = 0.015 (defective) ; q = 0.895 (not defective)
1. None of the antennas is defective.
P(200 not defective) = 0.895^200 = 2.315.10^-10
----

2. Three or more of the antennas are defective
P(3<= x <=200)
= 1 - P(0<= x <= 2)
= 1 - binomcdf(200,0.015,2)
= 0.5785
=================
Cheers,
Stan H